### v2

@@ 2.1. A Characterization of the Thought Experiment
In other words, for $\frac{3}{8}<\alpha\leq\frac{1}{2}$ and $\beta=\frac{1}{2}-\alpha$ the implied by Table 1 $\rho$ satisfies marginality but cannot be explained by the thought experiment. The violation of CHSH is maximal when $\alpha=\frac{1}{2}$ and $\beta=0$. That particular configuration was documented in Chambers et al. (2021).
@@
Next, we present an explanation of this unexpected behavior in the thought experiment and provide a necessary and sufficient condition for it. The condition is based on recent results from mathematical quantum physics that have studied a similar mathematical structure in other domains (Aubrun et al., 2021).
@@ 4.  A Necessary and Sufficient Condition for (Lack of) Entanglement of Choice
Chambers et al. (2021) provides a characterization of a version of our thought experiment with the random utility model (Block and Marschak, 1960, Falmagne, 1978, McFadden and Richter, 1990) restrictions on $A^{1,\Diamond}$ and $A^{2}$. They, however, only show the sufficiency of the unique representation property of $A^{1,\Diamond}$. Theorem 1 shows that in their setting, uniqueness is necessary while the random utility restriction is not. Li (2021) considers a setting similar to Chambers et al. (2021) but with finitely many DMs and at most three alternatives. As a result, the unique representation property of $A^{1,\Diamond}$ is satisfied. Because of that, the axiomatization in Li (2021) corresponds to the separable conditions generated by the interactions of the individual conditions. Our main result is a generalization of the characterizations in Chambers et al. (2021) and Li (2021).

### v3

@@ 2.1. A Characterization of the Thought Experiment
In other words, for $\frac{3}{8}<\alpha\leq\frac{1}{2}$ and $\beta=\frac{1}{2}-\alpha$ the implied by Table 1 $\rho$ satisfies marginality but cannot be explained by the thought experiment. The violation of CHSH is maximal when $\alpha=1/2$ and $\beta=0$. That particular configuration was documented in Chambers et al. (2021).
@@
Next, we describe a setting in which entangled choice arises. Consider the example in Table 1 corresponding to $\alpha=1/2$ and $\beta=0$. Sam and Frodo ate the big and the small cakes at a birthday celebration they had not been invited to. Eating a big cake is a severe crime in Hobbiton, while eating a small cake is a misdemeanor. After being caught, they can only be charged for one crime at a time. When either Frodo or Sam are facing the charge of eating the big cake they face the menu of actions $\{xw\}$: plead innocent ($x$) or plead guilty ($w$). When either Frodo or Sam are facing the charge of eating the small cake, the analogous actions $\{yz\}$ with $y$ meaning to plead innocent. With equal probabilities, they face either a soft or strict judge. They both get acquitted if they coordinate pleading innocent when the judge is soft. Similarly, they get a guilty plea deal when the judge is strict if they coordinate on pleading guilty. Frodo and Sam cannot communicate, but they can pay a lawyer to reveal to them the type of the judge they will face. Frodo is willing to pay the lawyer enough only if he faces the charge of eating the big cake. The same is true for Sam. When the lawyer gets paid enough by at least one of the them, she reveals the type of the judge to both of them truthfully. This means that they coordinate pleading at the same time innocent whenever the judge is soft and they both get acquitted. They plead guilty and get a guilty plea deal when the judge is strict. When they both face the charge of eating the small cake, they do not pay and the lawyer makes them anti-coordinate, maximizing their penalties. In this example, the joint behavior of Sam and Frodo is not separable. Entanglement arises because the randomization over actions is menu-dependent because of the information provided by a hidden agent, the lawyer.
Next, we explain this unexpected behavior in the thought experiment and provide a necessary and sufficient condition for it. The condition is based on recent results from mathematical quantum physics that have studied a similar mathematical structure in other domains (Aubrun et al., 2021).
@@ 4.  A Necessary and Sufficient Condition for (Lack of) Entanglement of Choice
Chambers et al. (2021) provides a characterization of a version of our thought experiment with the random utility model (Block and Marschak, 1960, Falmagne, 1978, McFadden and Richter, 1990) restrictions on $A^{1,\Diamond}$ and $A^{2}$. They, however, only show the sufficiency of the unique representation property of $A^{1,\Diamond}$. Theorem 1 shows that in their setting, uniqueness is necessary while the random utility restriction is not. Li (2021) considers a setting similar to Chambers et al. (2021) but with finitely many DMs and at most three alternatives. As a result, the unique representation property of $A^{1,\Diamond}$ is satisfied. Because of that, the axiomatization in Li (2021) corresponds to the separable conditions generated by the interactions of the individual conditions. Our main result is a generalization of the characterizations in Chambers et al. (2021) and Li (2021). Marginality is equivalent to a version of the separable choice model with negative probabilities (this result was shown for a general setup by Abramsky and Brandenburger, 2011 and for stochastic choice by Chambers et al., 2021).